# VDRive: Leveraging Reinforced VLA and Diffusion Policy for End-to-end Autonomous Driving

### v1

@@ 1 Introduction
However, the E2E paradigm relies on multimodal perception, fusion, and alignment (Hwang et al., 2024), where it aims to map raw sensor data directly to control actions through a unified learnable pipeline. A fundamental challenge lies in the significant representation gap between the high-dimensional, heterogeneous sensor data space (e.g., images, point clouds, audio) and the low-dimensional, structured action space (e.g., steering, throttle, braking). This gap is further exacerbated by semantic and temporal misalignment across modalities, as the features of the raw data operate at different levels of granularity (Li & Tang, 2024).
@@ 2.2 Diffusion Policy
Diffusion model-based approaches have proven their value in policy learning tasks (Chi et al., 2023; Yang et al., 2024). Since the advancement of diffusion models, they have gained recognition as a cornerstone in the field of generative modeling (Peebles & Xie, 2023). Conditioned diffusion models extend vanilla diffusion models by incorporating additional information during the generation process, while latent diffusion models improve computational efficiency and sample quality by operating in a compressed latent space (Rombach et al., 2022). Diffusion models have shown a strong potential to generate high-quality data in various modalities and to improve the representation of complex data structures (Yang & Wang, 2023). In robotics, Cheng Chi et al. (Chi et al., 2023), proposing a multimodal probabilistic action representation, considered the generation of robot action as a conditional diffusion denoising process. RDT-1B used a scalable Transformer backbone combined with diffusion models to capture the complexity and multimodality of bimanual actions, using diffusion models as a foundation model to effectively represent the multimodality inherent in bimanual manipulation tasks (Liu et al., 2024). In autonomous driving, the objectives of control policies are typically ill-defined and dynamically evolving rather than statically specified. The DIVER framework integrates diffusion models and reinforcement learning to address mode collapse in imitation learning for end-to-end autonomous driving. It proposes a Policy-Aware Diffusion Generator (PADG) that generates various trajectory candidates by conditioning on map elements, surrounding agents, and multiple reference ground truth trajectories (Song et al., 2025). As a planner, diffusion models excel in generating a wide distribution of diverse trajectories, while reinforcement learning can then optimize the generated trajectories based on specific diversity and safety rewards, effectively guiding the diffusion process to produce physically plausible and robust driving behaviors. Our framework jointly leverages diffusion policy and VLA via training data interaction and reinforcement learning, aiming at the efficient alignment of the sensor and action space.
@@ 5.2 Closed-loop Experiments
We evaluated VDRive using Dench2Drive, a closed-loop evaluation in the CARLA simulator. As shown in Table 2, our VDRive achieved a Driving Score of $66.25$, Success Rate of $50.51\%$ with Efficiency of $110.23$ and Comfortness of $22.90$. In Table 3, we report the Multi-ability results with Merging(22.35)(22.35), Overtaking(48.23)(48.23), Emergency Brake(50.96)(50.96), Give Way(40.00)(40.00), and Traffic Sign(66.72)(66.72).
@@ 5.4 Ablation Study
In the ablation study, we report the closed-loop results of different dataset constructions and different refinement module designs in the Bench2Drive-mini set. Table 4(a) shows that our proposed dataset construction that combines real-world, synthetic, and simulated scenes provides the best performance in closed-loop tasks. In Table 4(b), we report the results of different dynamic fusion approaches, where cross attention achieved the best performance compared to the refinement based on LSTM and GRU (Graves, 2012; Dey & Salem, 2017).

### v2

@@ 1 Introduction
However, the E2E paradigm relies on multimodal perception, fusion, and alignment (Hwang et al., 2024), where it aims to map raw sensor data directly to control actions through a unified learnable pipeline. A fundamental challenge lies in the significant representation gap between the high-dimensional, heterogeneous sensor data space (e.g., images, point clouds, audio) and the low-dimensional, structured action space (e.g., steering, throttle, braking). This gap is further exacerbated by semantic and temporal misalignment across modalities, as the features of the raw data operate at different levels of granularity (Li and Tang, 2024).
@@ 2.2 Diffusion Policy
Diffusion model-based approaches have proven their value in policy learning tasks (Chi et al., 2023; Yang et al., 2024). Since the advancement of diffusion models, they have gained recognition as a cornerstone in the field of generative modeling (Peebles and Xie, 2023). Conditioned diffusion models extend vanilla diffusion models by incorporating additional information during the generation process, while latent diffusion models improve computational efficiency and sample quality by operating in a compressed latent space (Rombach et al., 2022). Diffusion models have shown a strong potential to generate high-quality data in various modalities and to improve the representation of complex data structures (Yang and Wang, 2023). In robotics, Cheng Chi et al. (Chi et al., 2023), proposing a multimodal probabilistic action representation, considered the generation of robot action as a conditional diffusion denoising process. RDT-1B used a scalable Transformer backbone combined with diffusion models to capture the complexity and multimodality of bimanual actions, using diffusion models as a foundation model to effectively represent the multimodality inherent in bimanual manipulation tasks (Liu et al., 2024). In autonomous driving, the objectives of control policies are typically ill-defined and dynamically evolving rather than statically specified. The DIVER framework integrates diffusion models and reinforcement learning to address mode collapse in imitation learning for end-to-end autonomous driving. It proposes a Policy-Aware Diffusion Generator (PADG) that generates various trajectory candidates by conditioning on map elements, surrounding agents, and multiple reference ground truth trajectories (Song et al., 2025). As a planner, diffusion models excel in generating a wide distribution of diverse trajectories, while reinforcement learning can then optimize the generated trajectories based on specific diversity and safety rewards, effectively guiding the diffusion process to produce physically plausible and robust driving behaviors. Our framework jointly leverages diffusion policy and VLA via training data interaction and reinforcement learning, aiming at the efficient alignment of the sensor and action space.
@@ 5.2 Closed-loop Experiments
We evaluated VDRive using Bench2Drive, a closed-loop evaluation in the CARLA simulator. As shown in Table 2, our VDRive achieved a Driving Score of $66.25$, Success Rate of $50.51\%$ with Efficiency of $110.23$ and Comfortness of $22.90$. In Table 3, we report the Multi-ability results with Merging(22.35)(22.35), Overtaking(48.23)(48.23), Emergency Brake(50.96)(50.96), Give Way(40.00)(40.00), and Traffic Sign(66.72)(66.72).
@@ 5.4 Ablation Study
In the ablation study, we report the closed-loop results of different dataset constructions and different refinement module designs in the Bench2Drive-mini set. Table 4(a) shows that our proposed dataset construction that combines real-world, synthetic, and simulated scenes provides the best performance in closed-loop tasks. In Table 4(b), we report the results of different dynamic fusion approaches, where cross attention achieved the best performance compared to the refinement based on LSTM and GRU (Graves, 2012; Dey and Salem, 2017).